A unit cell of sodium chloride has four formula units. The edge length of the unit cell is 0.564 nm. What is the density of sodium chloride?   

Grade:upto college level

3 Answers

Deepak Patra
9 years ago
Hello Student,
Please find the answer to your question
Density of NaCI
= n * at. wt./Av. No. *a3 = 4 * 58.5/6.023 *1023 *(5.64 *10-8)3
= 2.16 g/cm3
Thanks
Deepak patra
askIITians Faculty
Ashutosh Tikoo
6 years ago
Using the formula ρ=N×a^3 / M × Avagadro numberWe getρ=4(0.564×10−7cm)3(58.5gmol−16.023×1023mol−1) =2.166gcm−3
Ashutosh Tikoo
6 years ago
Sorry I have written wrong formula rest of the answer is rightThe formula is ρ=(Number of atoms ×Molar mass=)/( a^3 × avagadro number)